OpenStudy (anonymous):

point-slope form y-b=m(x-a) (3,5) m=-1/2

OpenStudy (ksaimouli):

5-b=-1/2(3-a)

OpenStudy (anonymous):

how did you get that answer??

OpenStudy (ksaimouli):

 plug in m=-1/2

OpenStudy (ksaimouli):

Q?

OpenStudy (anonymous):

no, but thanks a lot